OpenStudy (anonymous):

rewrite in vertex form: y=-4x^2 -14x+8

OpenStudy (anonymous):

First we must know what vertex form is: y=a(x-h)^2 + k In order to rewrite in vertex form you need to complete the square: y= -4x^2 - 14x +8 You must first get rid of the -4 in front of the x^2, so you factor out the 4 as follows: y=-4(x^2- 14/4 x ) +8 Simplify y=-4(x^2-7/2x ) +8 Now we divide the term in front of x, which is -7/2, by 2 and square it and add it into the parenthesis. (-7/2)/2=-7/4 squaring it we get (-7/4)^2=49/16 y=-4(x^2 -7/2x + 49/16) + 8 But by doing the previous step we have the extra term -4* (49/16)=-49/4 so we add 49/4 at the end, y=-4(x^2- 7/2x + 49/16) + 8 + 49/4 We have completed completing the square we have y=-4(x- 7/4)^2 + (32+49)/4 So y=-4 (x- 7/4)^2 + 81/4 <=====ANSWER. Check: y= -4( x^2 -7/4 x -7/4 x +49/16)+81/4= -4(x^2 -14/4 x +49/16) + 81/4 = -4 (x^2 - 7/2 x +49/16) + 81/4 =-4x^2 -14x -49/4 +81/4 =-4x^2 -14x + 32/4 y=-4x - 14x+ 8 which is what we started with so our answer is correct.

OpenStudy (anonymous):

Sorry last line I forgot the " ^2" last line: y=-4x^2 -14x +8 Checks out :-)

OpenStudy (anonymous):

Thanks a lot!

OpenStudy (anonymous):

You are welcome uj1024 :-) Do you have any questions about any of the steps??? Please let me know. Otherwise have a good night (if you are in the US) :-)

OpenStudy (anonymous):

yeah im studying for finals next week...

OpenStudy (anonymous):

im glad i found this site